OpenStudy (anonymous):

im desperate for an answer to this and please answer and youll gain a fan: a length of #8 copper wire (radius = 1.63 mm) has a mass of 23.7 kg and a resistance of 2.061 ohm per km what is the over all resistance of the wire (d(copper) = 8.96 g/cm3 (cubed) i need the answer in three significant figures please help

OpenStudy (anonymous):

Volume will equal the cross sectional area times the length. you can find the length of the wire as follows: \[V=\pi r ^{2}l\] Solve this equation for l and you will get the length. Multiply this by the Ohm/km resistance for your answer. You will use the density as part of your unit conversion.

OpenStudy (anonymous):

hi hullsnipe, i have no idea what any of that means, im not in chemistry or anything im just trying to find the answer for my girlfriend or shes going to have to retake the whole class, i understand that this isnt freeanswers.com but if you could just solve it for me that would be great thank you

OpenStudy (anonymous):

\[\frac{ 27.3kg }{ 8.96\frac{ g }{ cm ^{3} } }\times \frac{ 1000g }{ kg }=3,046.875cm ^{3}=V\] To 3 sigfig the V is 3040 cm^3 Now: \[\frac{ 3040cm ^{3} }{(1.66mm)^{2} }\times \frac{ 100mm }{ cm }\times \frac{ 100mm }{ cm }\times \frac{ 1m }{ 100cm }\times \frac{ 1km }{1000 m }\times \frac{ 1 }{ 3.14 }=35.1km\] Now multiply this distance times the oms/km

OpenStudy (anonymous):

i have ADD i cant even read the first line without getting lost please sir if you know the answer just tell me, unlesss you're just trying to boost your smartscore by posting random things